A train travel 295km in 5hours.How many minutes will it covers in 8hours 15 minutes.
Answer by Alan3354(69443)

A train travel 295km in 5hours.
Find the speed.
d = r*t
r = d/t
--------------
How many minutes will it covers in 8hours 15 minutes.
15 minutes = 0.25 hours
d = r*t, r is the speed from above.